I typed the equation wrong!!!!!!!!!
This is what it should have been
2x -4y =8
x - 2y = 4
thanks,
dnn

Answer by funmath(2925)

Hi dnn,
L1) 2x -4y =8
L2) x - 2y = 4
:
This is the substitution method:
Solve L2) for x and subtitute that in for x in L1) and solve for y.
L2) x-2y=4
x-2y+2y=4+2y
x=4+2y <-- substitute this in for x in L1)
:
L1) 2(4+2y)-4y=8
8+4y-4y=8
8=8
Your varible was eliminated and 8 does = 8, the solution is all real numbers. This system is dependent, it's graphically the same line. Notice that if you were to divide all of the terms in L1 by 2, you would have L2.
Happy Calculating!!!